[Plate 1] FIGURE II-3: A false-color composite image of the merger sequence, with neutral hydrogen distribution shown in blue, the continuumsubtracted ionized hydrogen distribution shown in red, and the starlight from the R-band mosaic shown in yellowish-green and white. Bright stars that were saturated on the CCD appear red in this image. In the top row, from left to right: (a) Arp 295a&b; (b) NGC 4676a&b; and (c) NGC 520 and UCG 957. Bottom row, from the left, (d) NGC 3921; and (e) NGC 7252. (See page 37)

[Plate 2] FIGURE II-5: The intensity-weighted H i velocity field for each system, with the sense of velocity (red-shifted or blue-shifted with respect to the systemic velocity) indicated by the appropriate color. The systemic velocities appear either as green or yellow. In the following we list the heliocentric systemic velocities and velocity extrema, in km s^{−1}, after the system identification in the format (V_{min}, V_{sys}, V_{max}): Along the top row, from left to right: (a) Arp 295, (6500, 6860, 7150); (b) NGC 4676, (6300, 6600, 6900); (c) NGC 520, (2050, 2260, 2450). Along the bottom row are (d) NGC 3921, (5780, 5840, 5930); and (e) NGC 7252, (4600, 4740, 4880). The smooth runs of velocity support a tidal origin for the outer features. (See page 37)

FIGURE 3: Fate of original disk particles for equal mass encounters of varying inclinations, taken from figure 15 of Toomre & Toomre (1972). The letter “T” marks the location in the disk at pericenter for particles which are drawn into a tail, while a “B” denotes particles which get pulled into a bridge. Crosses indicate particles which become bound to the intruder, while filled circles indicate particles which escape. This figure shows that tails are pulled from the far side of the disk at orbital pericenter, while bridge material is pulled off of the near side.

FIGURE III-1: A composite image of the late-stage merger NGC 7252 showing: the summed B+R CCD mosaic image obtained using the CTIO 4 m (negative greyscales); the VLA C+D map of the integrated H i line emission (contours); and KPNO 2.1 m observations of the Hα line emission (white). North is up and east is to the left. The Hα data consist of three 4.′6 × 4.′6 CCD frames, one centered on the nucleus of NGC 7252 and one on the H ii region in each of the two tidal tails. The H i is located almost entirely in the outer tidal regions and extends beyond the end of the optical tails, particularly in the northwest.

FIGURE III-3: The H i line emission channel maps from the VLA C+D data running from 4536 km s^{−1} (bottom left panel) to 4877 km s^{−1} (top right panel). North is up and east is to the left. The two extreme crosses in each panel mark the locations of the H ii regions in the NW and E tidal tails while those near the middle mark the remnant center (left) and the brightest H ii region in the W loop (right). The solid contours are from the full resolution C+D data cube (FWHM: 26′′×16′′) and show column densities of (5, 15, 25, 35)×10^{19} cm^{−2}. The dotted contours from the smoothed C+D data (FWHM: 45′′×45′′) indicate a column density of 3×10^{19} cm^{−2}. The bottom three panels show emission from the W loop, with a large negative velocity relative to systemic (4740 km s^{−1}); the middle panels show predominantly the (blueshifted) NW tail; and the top panels show redshifted H i in the E tail and at the base of the NW tail.

[Plate 4] FIGURE IV-8: (a) The radial period versus pericentric distance for the tail particles, calculated according to their energies and angular momenta at T=72 (0.6 h^{−1} Gyr). R_{peri} is quantified in terms of R_{e} by the red horizontal dashed lines. The particles are color coded according to their return times and radii. Magenta denotes particles with a pericentric distance of less than 4R_{e} and a period shorter than the time at the end of the simulation (t_{end} =1.4 h^{−1} Gyr). Blue and green particles are those which are predicted to return to within 4R_{e} and 5R_{e} respectively between 1.4 and 3 h^{−1} Gyr. Red particles are expected to approach within 5R_{e} sometime between 3 h^{−1} Gyr and 10 h^{−1} Gyr of orbital periapse. Outer particles which will never fall to within 5R_{e} are drawn as black points. The large black dot is drawn at the location of the observed clump in the NW tail (HGvGS). The plot predicts that this clump will fall to within 14 h^{−1} kpc in 3.1 h^{−1} Gyr. (b) The sky view of the simulation at T=72, with the tail particles color coded as in (a). This figure is used to construct a “mask” used on the observational data to derived the quantities listed in Table 1. A large black dot is drawn in each panel at the location of the clump of gas and stars found in the NW tail (see HGvGS for details). Panel (a) predicts that this clump will have an orbital period of ∼4 h^{−1} Gyr and a pericenter of ∼14 h^{−1} kpc.

FIGURE AVIII-1: Starting geometry and orbital trajectories of each progenitor in the NGC 7252 simulation. (a) (left) As seen from above the orbital plane. The progenitor which gives rise to the NW tail is to the left while the E tail progenitor is to the right. (b) (right) As seen from the side. The progenitor to the NW tail is above and the E below.

TABLE II-4. Properties of Inner Regions

Line Quantity  units A295a A295b N4676a N4676b  N520^{(1)} N3921 N7252
(1) d_{opt}  h^{−1} kpc  34.2  17.9  14.0  14.0  21.8  15.5  14.6
(2) L_{R}  10^{10} h^{−2} L_{⊙}  1.2  1.0  0.8  0.9  0.9  1.5  2.0
(3) M_{HI}  10^{9} h^{−2} M_{⊙}  1.4  3.0  0.6  0.7 1.2 (1.8)  0.6  <0.1
(4) M_{gas}  10^{9} h^{−2} M_{⊙}  4.9  11.5  4.3  3.0 10.0 (10.6)  3.2  3.5
(5) M_{dyn}(d_{opt})^{(2)} 10^{10}h^{−1} M_{⊙}  30.0  10.5  5.2  9.1  7.7  11.9  4.0
(6) M_{H_{2}}/M_{HI}  α_{CO}  2.5  2.8  5.9  3.2 7.5 (4.9)  4.7  >70
(7) M_{gas}/L_{R}  M_{⊙}L_{⊙}^{−1}  0.4  1.1  0.5  0.3 1.1 (1.2)  0.2  0.2
(8) M_{gas}/M_{dyn}^{(3)}  h^{−1}  0.02  0.11^{∗}  0.08^{∗}  0.03  0.14^{∗}  0.03  0.09

Notes to TABLE II-4

(1) Parenthetical H i values for NGC 520 represent absorption-corrected quantities; see text.

(2) Dynamical mass calculated using d_{opt}, inclination and maximum velocity width in TABLE II-3, and prescription in Shostak 1978.

(3) An asterisk (^{∗}) denotes systems with observed minor-axis plumes of Hα.

TABLE II-5. Properties of Tidal Features

System:  Arp 295  NGC 4676  NGC 520  NGC 3921  NGC 7252
Feature:  S. Tail Bridge N. Tail S. Tail S. Tail Ring S. Tail N. Plume E. Tail NW. Tail
(1) l_{opt}  h^{−1} kpc  58  81  48  42  25  —  45  64  50  80
(2) l_{HI}  h^{−1} kpc  53  66  38  40  60  167  86  —  62  120
(3) L_{Hα+[NII]} 10^{40} h^{−2} erg s^{−1} 0.1  —  2.1  1.2  0.2  —  0.2  —  —  —
(4) L_{R}  10^{9} h^{−2} L_{⊙}  1.3  1.3  5.0  2.0  1.0  <0.1 2.0  3.7  0.6  1.7
(5) M_{HI}  10^{9} h^{−2} M_{⊙}  0.8  0.5  1.2  1.3  0.5 1.1  2.3  <0.05  0.6  1.2
(6) %L_{Hα+[NII]}  0.02  —  0.22 0.09  0.09 —  0.03  —  —  —
(7) %L_{R}  0.10 0.10  0.36 0.15  0.09 <0.01 0.09  0.16  0.02  0.05
(8) %M_{HI}  0.33 0.19  0.67 0.65  0.12 0.28 0.79  <0.02  0.30  0.60
(9) M_{HI}/L_{R}  M_{⊙}L_{⊙}^{−1}  0.6  0.4  0.2  0.6  0.5  >10  1.1  <0.02  1.0  0.7

Notes to TABLE II-5

(1)l_{opt} measured at 26.5 mag arcsec^{−2}.

(2)l_{HI} measured at 5×10^{19} cm^{−2}.

TABLE V-1. IR Selected Samples

(1)  (2)  (3)  (4)  (5)  (6)  (7)  (8)  (9)  (10)  (11)  (12)  (13)
Object  cz  log(L_{IR}) log(L_{K}) log(L_{B}) log(M_{H_{2}}) L_{IR}/L_{B} L_{IR}/L_{K} M_{H_{2}}/L_{B} M_{H_{2}}/L_{K} L_{IR}/M_{H_{2}} N_{tails} l_{tails}
(km s^{−1}) (L_{⊙})  (L_{⊙})  (L_{⊙})  (M_{⊙})  (M_{⊙}L_{⊙}^{−1}) (M_{⊙}L_{⊙}^{−1}) (L_{⊙}M_{⊙}^{−1})  (kpc)

IRAS BGS, log(L_{IR})>12

IR08572+3915 17,480  12.11  10.18 10.94  —  15  85  —  —  —  0  —
IR12112+0305 21,790  12.26  10.37 10.35  10.62  81  78  1.9  1.8  44  0  —
IR15250+3609 16,000  11.97  10.09 10.39  —  38  76  —  —  —  0  —
Arp 220  5,452  12.18  10.31 10.24  10.28  87  74  1.1  0.9  79  1?  33?
Mkn 273  11,300  12.14  10.56 10.58  10.23  36  38  0.4  0.5  81  1  29
IR05189-2524  12,820  12.09  —  10.47  10.36  41  —  0.8  —  54  0  —
IR22491-1808  23,170  12.09  —  10.55  10.43  34  —  0.8  —  46  2  40, 33
IR14348-1447  24,730  12.26  10.64 10.54  10.78  52  42  1.7  1.4  30  1  35
UGC 5101  12,000  12.00  10.66 10.42  10.08  38  22  0.5  0.3  83  1  39
Mkn 231  12,660  12.53  11.35 11.02  10.20  32  15  0.2  0.1  210  2?  27?, 22?

IRAS BGS, log(L_{IR})>11.5, Distance<100 Mpc

Arp 220  5,452  12.18  10.31 10.24  10.28  87  74  1.1  0.9  79  1?  33?
NGC 2623  5,545  11.55  10.02 10.33  9.77  17  34  0.3  0.6  60  2  28, 27
Arp 299  3,030  11.91  10.42 10.54  10.08  23  31  0.3  0.5  68  0  —
IC 883  7,000  11.60  10.12 10.37  9.87  17  30  0.3  0.6  54  1+1? 35, 30?
NGC 7469  4,963  11.60  10.29 10.63  9.96  9  20  0.2  0.5  44  0  —
NGC 1614  4,847  11.61  10.32 10.36  10.04  18  19  0.5  0.5  37  1  28
VV 114  6,028  11.62  10.34 10.45  10.71  15  19  1.8  2.3  8  0  —
NGC 6240  7,285  11.83  10.79 10.65  10.28  15  11  0.6  0.3  35  2?  28?, 27?

Notes to TABLE V-1

Distances and luminosities calculated using H_{o} = 75 km s^{−1} Mpc^{−1}

M_{H_{2}} is calculated assuming a conversion factor from integrated CO emissivity to H_{2} column density of 3×10^{20} molecules cm^{−2} “?” denote questionable or poorly delineated tidal tails.

TABLE V-2. Optically Selected Sample

(1)  (2)  (3)  (4)  (5)  (6)  (7)  (8)  (9)  (10)  (11)  (12) (13)
Object  cz  log(L_{IR}) log(L_{K}) log(L_{B}) log(M_{H_{2}}) L_{IR}/L_{B} L_{IR}/L_{K} M_{H_{2}}/L_{B} M_{H_{2}}/L_{K} L_{IR}/M_{H_{2}} N_{tails}  l_{tails}
(km s^{−1}) (L_{⊙}) (L_{⊙}) (L_{⊙}) (M_{⊙})  (M_{⊙}L_{⊙}^{−1}) (M_{⊙}L_{⊙}^{−1}) (L_{⊙}M_{⊙}^{−1})  (kpc)

Toomre Sequence of Ongoing Mergers from the NGC

NGC 4038/9 1,642 10.99 9.89 10.24 9.53  6  13  0.2  0.4  29  2  90, 50
NGC 4676  6,600 10.82 10.37 10.49 9.90  2  3  0.3  0.3  8  2  53, 50
NGC 7592  7,328 11.33 10.29 10.55 10.23  6  11  0.5  0.9  13  1+1? 90, 30
NGC 7764  1,706  —  —  9.87  —  —  —  —  —  —  2  40, 29
NGC 6621/2 6,284 11.26 10.53 10.60 10.00  5  5  0.3  0.3  18  1  70
NGC 3509  7,646 10.68  — 10.80  —  —  —  —  —  —  2  90, 33
NGC 520  2,160 10.88 10.06 10.21 10.08  5  7  0.7  1.0  6  1+1? 33, 27
NGC 2623  5,545 11.55 10.02 10.33 9.77  17  34  0.3  0.6  60  2  28, 27
NGC 3256  2,900 11.57 10.15 10.64 10.49  9  26  0.7  2.2  12  2  50, 60
NGC 3921  5,840 10.18 10.24 10.75 9.54  0.3  0.9  0.06  0.2  4  1  60
NGC 7252  4,750 10.70 10.33 10.81 9.67  0.8  2  0.07  0.2  11  2  93, 64

Notes to TABLE V-2

Distances and luminosities calculated using H_{o} = 75 km s^{−1} Mpc^{−1}

M_{H_{2}} is calculated assuming a conversion factor from integrated CO emissivity to H_{2} column density of 3×10^{20} molecules cm^{−2} “?” denote questionable or poorly delineated tidal tails.

FIGURE IV-8: [Plate 4](a) The radial period versus pericentric distance for the tail particles, calculated according to their energies and angular momenta at T=72 (0.6 h^{−1} Gyr). R_{peri} is quantified in terms of R_{e} by the red horizontal dashed lines. The particles are color coded according to their return times and radii. Magenta denotes particles with a pericentric distance of less than 4R_{e} and a period shorter than the time at the end of the simulation (t_{end} =1.4 h^{−1} Gyr). Blue and green particles are those which are predicted to return to within 4R_{e} and 5R_{e} respectively between 1.4 and 3 h^{−1} Gyr. Red particles are expected to approach within 5R_{e} sometime between 3 h^{−1} Gyr and 10 h^{−1} Gyr of orbital periapse. Outer particles which will never fall to within 5R_{e} are drawn as black points. (b) The sky view of the simulation at T=72, with the tail particles color coded as in (a). This figure is used to construct a “mask” used on the observational data to derived the quantities listed in Table 1. A large black dot is drawn in each panel at the location of the clump of gas and stars found in the NW tail (see HGvGS for details). Panel (a) predicts that this clump will have an orbital period of ∼4 h^{−1} Gyr and a pericenter of ∼14 h^{−1} kpc.